# msf-CNN: Patch-based Multi-Stage Fusion with Convolutional Neural Networks for TinyML

### v1

@@ 1 Introduction
  • We propose msf-CNN, a fusion-based approach to achieve ultra-low RAM footprint of neural network inference and we open-source its implementation;
@@
[…]
@@ 2 Background
[…]
@@ 3 High-Level Idea
As depicted in Figure 1(a), msf-CNN determines fusion settings (fusion position and depth), transforms layers accordingly into fusion blocks and rewrites global pooling and dense layers as their iterative implementation, which can further squeeze RAM usage without any computation overhead.
@@ 4 Problem Definitions & Assumptions
In Appendix B and Appendix C, we further detail the analysis of the cache buffer size and the amount of MAC operations.
@@ 5 DNN Graph Representation & Formulation
We interpret the optimization problems described in Section 4 by modeling the DNN as data-nodes graph. We transform the problem as a shortest path problem [31] and use off-the-shelf graph algorithms to find a solution that minimizes the peak memory usage as well as compute cost during inference regarding specified constraints.
@@ 5.1 DNN Representation
As described in Section 2, we model a DNN as a DAG $G=(V,E)$ with data nodes $v_{0},\dots,v_{n}$ representing input/output tensors of consecutive layers and $m$ edges $e_{1},\dots,e_{m}$ that represent single layers or fusion blocks. Each edge is also encoded with resource requirements by layer or fusion block. Specifically, the first ($v_{0}$) and the last node ($v_{n}$) are the input and output tensor of the neural network, respectively.
@@
A typical example depicted in Figure 1(b) explains how to use DAG for representing a simple neural network. Tensors are transformed into nodes, operators and fusion blocks are edges. Edges are encoded with RAM usages and MAC amounts of their corresponding operators. Hence, the problem is transformed to find an optimal path from the input node to the output node of the graph.
@@ 5.2 Encoding RAM Usage
where $I$ and $O$ are the size of input and output tensor, respectively. $Buf$ represents the cache buffer size of the fusion block, which is determined by the chosen cache scheme. In this work it is given in Equation 11. Trivially, for non-fused layers $Buf$ is always set to zero since no fusion cache is needed.
@@ 5.3 Encoding Compute Cost
The encoding steps of compute cost are similar to encoding peak memory usage. Here we use MAC operations as the indicator of compute cost. In this paper, the MAC amount of fusion block is given in Equation 14 and Equation 15.
@@
Therefore, the compute overhead factor $F$ representing the ratio of the MAC amount after fusion to the vanilla, common one without fusion is expressed as $F=C_{S}/C_{vanilla}$. For the constraints in Equation 2, users can set a maximum compute overhead factor $F_{max}$ expressed as $F_{max}=C_{max}/C_{vanilla}$. In the following sections, we will discuss several graph-based algorithms to solve the optimization problem.
@@ 6.1 Problem P1: Minimizing Peak RAM Usage
The unconstrained optimization is to find a complete compute path with minimal peak RAM usage, which is equivalent to finding the path that minimizes the maximum weight of edges (minimax path problem). As mentioned above, this can be solved by modified Dijkstra algorithm. An example path with minimal peak RAM usage is presented in Figure 1(b).
For the constraint of compute cost limit (Equation 2), the pruning strategy needs co-design with its optimization problem (Equation 1). We noticed that all possible peak RAM usages have already been encoded into the edges. Therefore, the problem can be transformed into the following: we first construct a candidate solution set $\{S_{0},S_{1},\dots,S_{i},\dots\}$ with
@@
all edges in Gi1 with the maximal RAM $\text{all edges in $G_{i-1}$ with the maximal RAM usage},$ (9)
@@ 6.2 Problem P2: Minimizing Compute Cost
We first discuss the unconstrained variant, which is identical to $P_{max}=\infty$. In this case, finding the solution is equivalent to finding the shortest complete compute path – the path with a minimal sum of MAC – of the graph, which can be again solved by classical algorithm like Dijkstra’s [7]. Figure 1(b) shows an example with an optimal path marked in red.
@@ 6.3 Analytical Results
As shown in Table 1, both optimizers can indeed theoretically suppress the peak RAM usage without violating all preset constraints. The high RAM usage compression is achieved with increase of deep fusion blocks, thereby introducing a high compute overhead. The extreme cases lay on the unconstrained optimization minimizing the RAM usage by more than $90\%$, while reluctantly introducing 1.6×1.6$\times$1.6 × to 2.7×2.7$\times$2.7 × of compute overhead. This is only suitable for time-intensive applications with a high limited RAM budget.
@@
[…]
We also conducted a preliminary analysis of the heuristic strategy used in MCUNetV2, which fuses only the early layers to minimize RAM usage. While this approach is simple and straightforward, it tends to yield suboptimal fusion configurations. As shown in Table 1, msf-CNN discovers better solutions and offers greater flexibility compared to the heuristic strategy. The analytical results were further validated by on-board experiments presented in Section 8.
[…]
@@
We have implemented the msf-CNN fusion mechanism on top of microTVM v0.16.0 [6]. We use the TVM frontend to convert models into intermediate representation (IR), and rewrite the compute graph and low-level routines of operators to fit the fusion settings. We leveraged RIOT-ML [14] to benchmark the fused models (transform into C code by microTVM) on the IoT boards shown in Table 4.
@@
Iterative Computation of Global Pooling – As illustrated in Figure 3, standard global pooling requires that all elements of the input tensor stored in RAM. In our approach, the global pooling layer receives one or a few input elements at each step and iteratively updates the result. For a $7\times 7$ global pooling layer, this allows us to compress the RAM usage to 2% of the original size, without introducing any redundant computations or computation overhead.
Iterative Computation of Dense Layer – We noted that the matrix multiplication in dense layers can be implemented by splitting the input vector into individual elements, multiplying each element with its corresponding weight column, and iteratively summing the results, as shown in Figure 3. Unlike the original approach, which requires the complete input tensor, this method processes only one element of the input tensor per iteration. For a 1024→256 dense layer, this approach compresses memory usage to 20% of the original.
@@ 8 Experiments on Microcontrollers
More concretely, we measured peak RAM usage and compute latency based on the fusion settings in Section 6.3, as reported in the following. As shown in Appendix E, we carried out our experiments on the relevant 32-bit microcontroller architectures: Arm Cortex-M, Espressif Xtensa, and RISC-V. For our model zoo, we chose MBV2-w0.35, MN2-vww5 and MN2-320K as they are good representatives of backbones for applications in AIoT [29], as also used in prior works [22, 40]. We compare msf-CNN performance to the closest related work: MCUNetV2 [22] and StreamNet-2D [40], more simply denoted StreamNet in the following.
@@ 8.1 Minimal Peak RAM Usage
First, we evaluated solutions to P1 while relaxing Equation 2, i.e. the fusion settings with minimum peak RAM usage, without considering time constraint. Results are shown in Table 3. We observe that, compared to prior art (StreamNet-2D and MCUNetV2), msf-CNN can further reduce the peak RAM usage by 65% to 87%. We could even deploy the MBV2-w0.35 model onto the SiFive board that provides only 16 kB RAM (!). However, achieving this high compression ratio comes at the expense of increased computational latency, which we measured in Table 3. Interestingly, while clock frequency plays a decisive role, MCU architecture can also have a crucial effect, for larger models. For instance, notice latency with Xtensa esp32s3 at 240MHz versus RISC-V esp32c3 at 160 MHz, for the MN2-320K model (in Table 3). Nevertheless, we measured that latency increases 2×2$\times$2 × to 5×5$\times$5 × compared to vanilla (non-fused) CNN. Hence, such minimal RAM settings are only suitable for latency-tolerant applications on the smallest devices.
@@ 8.2 Impact of RAM Budget Limit
As shown in Figure 4 and Table 5, the measured peak RAM usage consistently obeys to the given constraints, thereby validating the correctness of the optimizer and corroborating our analytical results. Based on these, we observe that higher RAM budgets result in shorter compute latency for the optimal fusion configurations identified by msf-CNN. This is because the optimizer tends to favor configurations with either no fusion or shallow fusion depths, which correspond to higher peak RAM usage but lower computational costs.
@@
[…]
@@ 8.3 Impact of Computation Cost Limit
Our experiments demonstrate msf-CNN’s capability to optimize resource usage with diverse CNN models, under user-specified constraints emphasizing either compute latency or RAM footprint. Furthermore, msf-CNN generates code that is deployable across diverse MCUs’ ISAs. Users can thus produce optimal CNN fusion configurations tailored to specific industrial hardware requirements. However, some limitations (Appendix F) remain, on which our future work will focus next.
@@ 9 Related Work
Tiny Neural Architecture Search (NAS) – This technique is employed to automatically search for model structures with optimal accuracy under the constraints of memory, flash footprint and compute latency. TinyNAS [21], $\mu$NAS [20] and the Once-for-All Network [5] leverage Neural Architecture Search (NAS) to design CNNs with exceptionally small memory requirements for MCUs. The resulting networks require only a few hundred kilobytes of RAM for execution. However, contrary to msf-CNN, these methods necessitate retraining or fine-tuning of pre-existing networks.
@@ 10 Conclusion
Convolutional neural networks (CNNs) must not only execute in the cloud or on edge computing gateways, but also on the smaller, more energy-efficient microcontroller-based devices which take part in our cyber-physical systems. Microcontrollers pose a great challenge for CNNs regarding the joint optimization of RAM memory consumption and inference latency. In this context, we presented msf-CNN, a technique and heuristics able to identify pools of practical patch-based fusion optimizations for CNN inference, which jointly satisfy memory and latency constraints. Compared to previous work on CNN fusion for microcontrollers, msf-CNN identifies a wider set of applicable solutions, on much more diverse hardware. Our experimental evaluation using the open source implementation we provide for common microcontrollers (ARM Cortex-M, RISC-V, and ESP32) show that msf-CNN can achieve inference with less than 50% the peak RAM usage compared to state-of-the-art. As such, msf-CNN provides a new level of flexibility for embedded system designers, which can now better tune the trade-off between peak RAM and model inference latency on various MCUs.
@@ Appendix D Complexity analysis of the search algorithm
Appendix E Experiment Details
Here, we provide additional details of the experiments, including hardware specifications and results not presented in the main text.
E.1 Hardware Settings
Table 4 presents the evaluation board and microcontrollers used in this study, along with detailed specifications including RAM size, flash memory capacity, and CPU characteristics.
Board MCU Core RAM Flash
Nucleo-f767zi STM32F767ZI Cortex-M7 @ 216 MHz 512 2048
Stm32f746g-disco STM32F746NG Cortex-M7 @ 216 MHz 320 1024
Nucleo-f412zg STM32F412ZG Cortex-M4 @ 100 MHz 256 1024
esp32s3-devkit ESP32-S3-WROOM-1N8 Xtensa @ 240 MHz 512 8192
esp32c3-devkit ESP32c3-1-MINI-M4N4 RISC-V @ 160 MHz 384 4096
hifive1b SiFive FE310-G002 RISC-V @ 320 MHz 16 4096
E.2 Supplemental Results
Table 5 provides the underlying data corresponding to Figure 4, including the specific constraints applied to P1 and P2, respectively.
Table 5: Optimal fusion settings on Nucleo-f767zi. RAM (kB), Latency (ms). SAA: Same as above. Bold: msf-CNN beats MCUNetV2.

### v3

@@ 1 Introduction
  • We propose msf-CNN, a fusion-based approach to achieve ultra-low RAM footprint of neural network inference and we open-source its implementation [^1] Please check https://github.com/TinyPART/msf-CNN;
@@
  • We designed a pruning strategy to squeeze the search space and use graph-based algorithm to find solutions in reasonable time complexity (from $O(2^{N-2})$ to polynomial time).
@@
While our main focus is on microcontrollers, msf-CNN is not limited to them. Its cost estimator and C-code backend also support general CPU platforms (e.g., x86, Cortex-A), enabling broader use such as memory-optimized cloud inference. Moreover, with appropriate cost models and backends, msf-CNN can be extended to accelerators (GPUs, FPGAs, ASICs). We leave non-MCU experiments to future work to maintain focus on devices with tighter RAM constraints.
@@ 2 Background
Patch-based Fusion for DNN on FPGA & GPU – Patch-based fusion was initially proposed in [2] as a fusion scheme for Convolutional Neural Network (CNN) deployed on Field Programmable Gate Array (FPGA) to reduce the off-chip DRAM usage and I/O overhead. Instead of computing the complete feature maps for each layer, it fuses convolutional layers into a single block (pyramid structure) and computes only one or a few output elements. This approach requires only small portions (tiles) of the feature maps loaded onto DRAM. However, the reduction of RAM is at the cost of re-computing the overlapped elements in feature maps required by adjacent fused layers. DeFiNES [28], another fusion framework, explored different cache strategies within fused layers to alleviate the re-computation issue. (Fully-recompute, H-Cached & V-recompute, and Fully-cache). Fully-recompute eliminates caching entirely, requiring all overlapping input tensor elements to be recalculated; H-cached & V-recompute caches elements along the horizontal axis while recomputing vertical overlaps; and Fully-cache retains all overlapping elements in memory. These approaches illustrate a critical trade-off—enhanced caching progressively reduces compute redundancy but proportionally increases RAM usage, with cached element quantity inversely correlating to compute overhead and directly scaling with memory demands. Additional work has also applied fusion on GPUs, for instance [31] used it for cancer detection in medical pictures. Note that patch-based fusion is fundamentally different from kernel fusion techniques. We elaborate on this distinction and its implications in Appendix˜A.
@@ 3 High-Level Idea
As depicted in Figure˜1(a), msf-CNN determines fusion settings (fusion position and depth), transforms layers accordingly into fusion blocks and rewrites global pooling and dense layers as their iterative implementation, which can further squeeze RAM usage without any computation overhead.
@@ 4 Problem Definitions & Assumptions
In Appendix˜B and Appendix˜C, we further detail the analysis of the cache buffer size and the amount of MAC operations.
@@ 5 DNN Graph Representation & Formulation
We interpret the optimization problems described in Section˜4 by modeling the DNN as data-nodes graph. We transform the problem as a shortest path problem [35] and use off-the-shelf graph algorithms to find a solution that minimizes the peak memory usage as well as compute cost during inference regarding specified constraints.
@@ 5.1 DNN Representation
As described in Section˜2, we model a DNN as a DAG $G=(V,E)$ with data nodes $v_{0},\dots,v_{n}$ representing input/output tensors of consecutive layers and $m$ edges $e_{1},\dots,e_{m}$ that represent single layers or fusion blocks. Each edge is also encoded with resource requirements by layer or fusion block. Specifically, the first ($v_{0}$) and the last node ($v_{n}$) are the input and output tensor of the neural network, respectively.
@@
A typical example depicted in Figure˜1(b) explains how to use DAG for representing a simple neural network. Tensors are transformed into nodes, operators and fusion blocks are edges. Edges are encoded with RAM usages and MAC amounts of their corresponding operators. Hence, the problem is transformed to find an optimal path from the input node to the output node of the graph.
@@ 5.2 Encoding RAM Usage
where $I$ and $O$ are the size of input and output tensor, respectively. $Buf$ represents the cache buffer size of the fusion block, which is determined by the chosen cache scheme. In this work it is given in Equation˜11. Trivially, for non-fused layers $Buf$ is always set to zero since no fusion cache is needed.
@@ 5.3 Encoding Compute Cost
The encoding steps of compute cost are similar to encoding peak memory usage. Here we use MAC operations as the indicator of compute cost. In this paper, the MAC amount of fusion block is given in Equation˜14 and Equation˜15.
@@
Therefore, the compute overhead factor $F$ representing the ratio of the MAC amount after fusion to the vanilla, common one without fusion is expressed as $F=C_{S}/C_{vanilla}$. For the constraints in Equation˜2, users can set a maximum compute overhead factor $F_{max}$ expressed as $F_{max}=C_{max}/C_{vanilla}$. In the following sections, we will discuss several graph-based algorithms to solve the optimization problem.
@@ 6.1 Problem P1: Minimizing Peak RAM Usage
The unconstrained optimization is to find a complete compute path with minimal peak RAM usage, which is equivalent to finding the path that minimizes the maximum weight of edges (minimax path problem). As mentioned above, this can be solved by modified Dijkstra algorithm. An example path with minimal peak RAM usage is presented in Figure˜1(b).
For the constraint of compute cost limit (Equation˜2), the pruning strategy needs co-design with its optimization problem (Equation˜1). We noticed that all possible peak RAM usages have already been encoded into the edges. Therefore, the problem can be transformed into the following: we first construct a candidate solution set $\{S_{0},S_{1},\dots,S_{i},\dots\}$ with
@@
[…]
@@ 6.2 Problem P2: Minimizing Compute Cost
We first discuss the unconstrained variant, which is identical to $P_{max}=\infty$. In this case, finding the solution is equivalent to finding the shortest complete compute path – the path with a minimal sum of MAC – of the graph, which can be again solved by classical algorithm like Dijkstra’s [8]. Figure˜1(b) shows an example with an optimal path marked in red.
@@ 6.3 Analytical Results
As shown in Table˜1, both optimizers can indeed theoretically suppress the peak RAM usage without violating all preset constraints. The high RAM usage compression is achieved with increase of deep fusion blocks, thereby introducing a high compute overhead. The extreme cases lay on the unconstrained optimization minimizing the RAM usage by more than $90\%$, while reluctantly introducing $1.6\times$ to $2.7\times$ of compute overhead. This is only suitable for time-intensive applications with a high limited RAM budget.
@@
We also conducted a preliminary analysis of the heuristic strategy used in MCUNetV2, which fuses only the early layers to minimize RAM usage. While this approach is simple and straightforward, it tends to yield suboptimal fusion configurations. As shown in Table˜1, msf-CNN discovers better solutions and offers greater flexibility compared to the heuristic strategy. The analytical results were further validated by on-board experiments presented in Section˜8.
Table 1: Analytical results with msf-CNN under different constraints. Vanilla: un-fused models. Heuristic: minimize RAM consumption by only fusing heading layers. SAA: Same as above. Gray: msf-CNN beats heuristic.
@@
We have implemented the msf-CNN fusion mechanism on top of microTVM v0.16.0 [7]. We use the TVM frontend to convert models into intermediate representation (IR), and rewrite the compute graph and low-level routines of operators to fit the fusion settings. We used RIOT-ML [17] to benchmark the fused models: the models are transformed into C code using microTVM and integrated in dedicated firmware leveraging a common IoT operating system (RIOT [4]) which we can run on various boards shown in Table˜2.
@@
Iterative Computation of Global Pooling – As illustrated in Figure˜3, standard global pooling requires that all elements of the input tensor stored in RAM. In our approach, the global pooling layer receives one or a few input elements at each step and iteratively updates the result. For a $7\times 7$ global pooling layer, this allows us to compress the RAM usage to 2% of the original size, without introducing any redundant computations or computation overhead.
Iterative Computation of Dense Layer – We noted that the matrix multiplication in dense layers can be implemented by splitting the input vector into individual elements, multiplying each element with its corresponding weight column, and iteratively summing the results, as shown in Figure˜3. Unlike the original approach, which requires the complete input tensor, this method processes only one element of the input tensor per iteration. For a 1024→256 dense layer, this approach compresses memory usage to 20% of the original.
@@ 8 Experiments on Microcontrollers
More concretely, we measured peak RAM usage and compute latency based on the fusion settings in Section˜6.3, as reported in the following. As shown in Table˜2, we carried out our experiments on the relevant 32-bit microcontroller architectures: Arm Cortex-M, Espressif Xtensa, and RISC-V. For our model zoo, we chose MBV2-w0.35, MN2-vww5 and MN2-320K as they are good representatives of backbones for applications in AIoT [33], as also used in prior works [26, 46]. We compare msf-CNN performance to the closest related work: MCUNetV2 [26] and StreamNet-2D [46], more simply denoted StreamNet in the following.
Board MCU Core RAM Flash
Nucleo-f767zi STM32F767ZI Cortex-M7 @ 216 MHz 512 2048
Stm32f746g-disco STM32F746NG Cortex-M7 @ 216 MHz 320 1024
Nucleo-f412zg STM32F412ZG Cortex-M4 @ 100 MHz 256 1024
esp32s3-devkit ESP32-S3-WROOM-1N8 Xtensa @ 240 MHz 512 8192
esp32c3-devkit ESP32c3-1-MINI-M4N4 RISC-V @ 160 MHz 384 4096
hifive1b SiFive FE310-G002 RISC-V @ 320 MHz 16 4096
@@ 8.1 Minimal Peak RAM Usage
First, we evaluated solutions to P1 while relaxing Equation˜2, i.e. the fusion settings with minimum peak RAM usage, without considering time constraint. Results are shown in Table˜4. We observe that, compared to prior art (StreamNet-2D and MCUNetV2), msf-CNN can further reduce the peak RAM usage by 65% to 87%. We could even deploy the MBV2-w0.35 model onto the SiFive board that provides only 16 kB RAM (!). However, achieving this high compression ratio comes at the expense of increased computational latency, which we measured in Table˜4. Interestingly, while clock frequency plays a decisive role, MCU architecture can also have a crucial effect, for larger models. For instance, notice latency with Xtensa esp32s3 at 240MHz versus RISC-V esp32c3 at 160 MHz, for the MN2-320K model (in Table˜4). Nevertheless, we measured that latency increases $2\times$ to $5\times$ compared to vanilla (non-fused) CNN. Hence, such minimal RAM settings are only suitable for latency-tolerant applications on the smallest devices.
@@ 8.2 Impact of RAM Budget Limit
As shown in Figure˜4 and Table˜6, the measured peak RAM usage consistently obeys to the given constraints, thereby validating the correctness of the optimizer and corroborating our analytical results. Based on these, we observe that higher RAM budgets result in shorter compute latency for the optimal fusion configurations identified by msf-CNN. This is because the optimizer tends to favor configurations with either no fusion or shallow fusion depths, which correspond to higher peak RAM usage but lower computational costs.
@@
Figure 4: Trade-off between RAM and latency of different optimal fusion settings on Nucleo-f767zi. P1: Minimize RAM s.t. compute cost limit. P2: Minimize compute cost s.t. RAM limit. For more detailed results, please refer to Table˜6 in Appendix˜F.
@@ 8.3 Impact of Computation Cost Limit
msf-CNN underperforms StreamNet because our implementation does not use hardware-dependent libraries or acceleration instructions, while StreamNet is optimized for ARM platforms via the CMSIS [22] and employs a 2-D cache (ours uses 1-D), which further mitigates recomputation. To enable a fairer comparison, we added an optional CMSIS backend to msf-CNN for ARM devices and conducted additional measurements, as shown in Figure˜4. The results show msf-CNN’s Pareto curves of CMSIS variants are much closer to StreamNet’s. We are also extending msf-CNN with 2-D cache support to provide greater flexibility across hardware platforms.
9 Discussion
Our experiments demonstrate msf-CNN’s capability to optimize resource usage with diverse CNN models, under user-specified constraints emphasizing either compute latency or RAM footprint. Furthermore, msf-CNN generates code that is deployable across diverse MCUs’ ISAs. Users can thus produce optimal CNN fusion configurations tailored to specific industrial hardware requirements. However, some limitations remain, on which our future work will focus next:
Neural Network Architecture – The work currently focuses exclusively on convolutional neural network architectures (CNNs). The analysis of other prevalent structures, particularly attention mechanisms and recurrent neural networks (RNNs), remains an open research direction. Please check Appendix˜G for the current state of extension.
@@ 10 Related Work
Memory Optimization for CNN layers – Memory optimization strategies can be broadly categorized into scheduling-based and fusion-based methods. Scheduling-based methods, such as those implemented in frameworks like Ansor [47], vMCU [49], MoDEL [37] and TinyEngine [26], focus on the efficient reuse of memory pools to minimize peak memory usage by leveraging the different lifetimes of inter- and intra-layer tensors. Although both methods achieve a peak memory reduction exceeding 50%, they still generate a complete output tensor for each layer. This requirement remains problematic for low-power MCUs with limited RAM. Prior work on fusion was covered in Section 2. Contrary to msf-CNN, these methods do not fully exploit the potential of multiple fusion blocks.
@@ 11 Conclusion
Convolutional neural networks (CNNs) must not only execute in the cloud or on edge computing gateways, but also on the smaller, more energy-efficient microcontroller-based devices which take part in our cyber-physical systems. Microcontrollers pose a great challenge for CNNs regarding the joint optimization of RAM memory consumption and inference latency. In this context, we presented msf-CNN, a technique and heuristics able to identify pools of practical patch-based fusion optimizations for CNN inference, which jointly satisfy memory and latency constraints. Compared to previous work on CNN fusion for microcontrollers, msf-CNN identifies a wider set of applicable solutions, on much more diverse hardware. Our experimental evaluation using the open source implementation we provide for common microcontrollers (ARM Cortex-M, RISC-V, and ESP32) show that msf-CNN can achieve inference with less than 50% the peak RAM usage compared to state-of-the-art. As such, msf-CNN provides a new level of flexibility for embedded system designers, which can now better tune the trade-off between peak RAM and model inference latency on various hardware.
Acknowledgment
The authors thank Adrien Tousnakhoff for useful discussions. Work contributing to these results was partially funded by the ANR France 2030 Programme (ANR-22-PTCC-0001 and ANR-22-PEFT-0007).
@@ Appendix D Complexity analysis of the search algorithm
Appendix E Accuracy Evaluation
msf-CNN is a computation scheduling and memory optimization technique that does not alter the model’s architecture, parameters, or the mathematical operations performed. Like MCUNet and StreamNet, msf-CNN only changes when and how intermediate results are computed and stored to minimize peak memory. Therefore, the final output, and consequently the model’s accuracy, remains identical to the original, unfused model. Hence, standard ML performance benchmarks focusing on accuracy are irrelevant here.
Nevertheless, to make sure, we conducted extra experiments on imagenet and vww dataset, as MCUNet did. We reused the model weights from the pre-trained MCUNet models, and selected $P_{max}=64kB$ as fusion constraints. The results show in Table˜5 that no Top-1 accuracy drop was found between vanilla models and the msf-CNN variants.
Table 5: Model Top-1 Accuracy
Method MBV2-w0.35 MN2-vww5 MN2-320K
Vanilla 48.94 % 88.89 % 61.76 %
msf-CNN 48.94 % 88.89 % 61.76 %
Appendix F Experiment Details
Here, we provide additional details of the experiments, including results not presented in the main text.
F.1 Supplemental Results
Table˜6 provides the underlying data corresponding to Figure˜4, including the specific constraints applied to P1 and P2, respectively.
Table 6: Optimal fusion settings on Nucleo-f767zi. RAM (kB), Latency (ms). SAA: Same as above. Gray: msf-CNN beats MCUNetV2.